Two cars leave an intersection one car travels north; the other east. When the car traveling north had gone 6 miles, the distance between the cars was 2 miles more than the distance traveled by the car heading east how far had the eastbound car traveled
Answer by stanbon(75887)

Two cars leave an intersection one car travels north; the other east. When the car traveling north had gone 6 miles, the distance between the cars was 2 miles more than the distance traveled by the car heading east how far had the eastbound car traveled
---
Draw the picture.
north:: 6 miles
east:: x miles
diagonal(x+2) miles
---------
Equation:
(x+2)^2 = x^2 + 6^2
-----
x^2 + 4x + 4 = x^2 + 36
------
4x - 32 = 0
4(x-8) = 0
x = 8 miles
x+2 = 10 miles
----------------
Cheers,
Stan H.
------------